# ConsistencyDet: Robust Object Detector with Denoising Paradigm of Consistency Model

### v1

@@ 2.4 Nomenclature
3 Proposed detection method
3.1 Preliminaries
@@
$p_{\theta}^{(t)}(x_{t-1}|x_{t})=\begin{cases}\mathcal{N}(f_{\theta}^{(1)}(x_{1}),\sigma_{1}^{2}\mathbf{I})&\text{if }t=1,\\ q_{\sigma}(x_{t-1}|x_{t},f_{\theta}^{(t)}(x_{t}))&\text{otherwise}.\end{cases}$ (2)
@@
Consistency Model. In this study, within the framework of the Consistency Model that utilizes deep neural networks, we investigate two cost-effective methodologies for enforcing boundary conditions. Let $F_{\theta}(x,t)$ represent a free-form deep neural network whose dimensionality is same as $x$. The first method directly parameterizes the Consistency Model as:
$f_{\theta}(x,t)=\begin{cases}x&\text{if }t=\tau,\\ F_{\theta}(x,t)&\text{if }t\in[\tau,T),\end{cases}$ (4)
@@
where $\tau$ is an integer within the range $[0,T-1]$. The second method parameterizes the Consistency Model by incorporating skip connections and is formalized as follows:
$f_{\theta}(x,t)=c_{\text{skip}}(t)x+c_{\text{out}}(t)F_{\theta}(x,t),$ (5)
@@
where $c_{\text{skip}}(t)$ and $c_{\text{out}}(t)$ are differentiable functions [22], satisfying $c_{\text{skip}}(\tau)=1$ and $c_{\text{out}}(\tau)=0$. By employing this construction, the Consistency Model becomes differentiable at $t=\tau$, provided that $F_{\theta}(x,t)$, $c_{\text{skip}}(t)$, and $c_{\text{out}}(t)$ are all differentiable. This differentiability is crucial for the training of continuous-time Consistency Models.
@@
Figure 3: Training procedures of the proposed ConsistencyDet. After extracting features through the backbone, random Gaussian noise is added to GT boxes following the Consistency Model’s noise addition strategy. These noised boxes with corresponding features processed by RoI pooler are then input to ConsistencyHead for iterative noise removal, with several basic modules, ultimately yielding the final detection results. Each basic module contains self-attention mechanism (SA), dynamic convolutional layers (DC) and the head (HD) of classification and regression.
@@
Figure 4: To ensure the self-consistency property of Consistency Model, it requires feeding the noised boxes corresponding to $(t-1)$-th and $t$-th time steps into the model simultaneously. They jointly predict the final results, then compare them with GT to estimate the training loss.
Training process of Consistency Model. Within the context of the Consistency Model, this study introduces two distinct training strategies that capitalize on the self-consistency property. The proposed strategies encompass distillation training, which transfers knowledge from a Diffusion Model to the Consistency Model, and an autonomous training regime wherein the Consistency Model is trained independently.
@@
The first method employs numerical ODE solvers in conjunction with a pre-trained diffusion-based detector, referred to as DiffusionDet, to generate pairs of proximate points along a PF ODE trajectory. By minimizing the divergence between the model predictions for these point pairs, knowledge from the Diffusion Model is effectively transferred to the Consistency Model. As a result, this allows for the generation of high-quality samples through a single evaluation of the network [35, 36]. The comprehensive training protocol for the Consistency Model through distillation is encapsulated in Algorithm 1.
Conversely, the second approach dispenses with the need for a pre-trained Diffusion Model, thereby enabling the Consistency Model to be trained in a standalone manner. This technique delineates the Consistency Model as an independent entity within the broader spectrum of its applications. The detailed training process for the independently trained Consistency Model is encapsulated in Algorithm 2.
@@
1 $\theta^{-}\leftarrow\theta$ while not convergence do
@@ 7       $\theta^{-}\leftarrow\text{stopgrad}(\mu\theta^{-}+(1-\mu)\theta)$;
1 $\theta^{-}\leftarrow\theta$ and $k\leftarrow 0$ while not convergence do
2       Sample $\mathbf{x}\in\mathcal{D}$, and $t\in\mathcal{U}\left[1,T(k)-1\right]$; Sample $\epsilon\in\mathcal{N}(0,I)$; $\mathcal{L}(\theta,\theta^{-})\leftarrow\lambda(\sigma_{t})d(f_{\theta}\left(\mathbf{x}+\epsilon\sigma_{t+1},\sigma_{t+1}\right),f_{\theta^{-}}\left(\mathbf{x}+\epsilon\sigma_{t},\sigma_{t}\right))$; $\theta\leftarrow\theta-\eta\nabla_{\theta}\mathcal{L}(\theta,\theta^{-})$; $\theta^{-}\leftarrow\text{stopgrad}(\mu(k)\theta^{-}+(1-\mu(k))\theta)$; $k\leftarrow k+1$;
3 end while
@@ 3.2 Architecture
ConsistencyDet is composed of an image encoder and a detection decoder, with the workflow depicted in Fig. 3. The input to the image encoder is the original image along with its GT boxes. These GT boxes are then subjected to a noise injection procedure as part of the Consistency Model, where Gaussian noise is randomly introduced. The perturbed boxes are processed by the RoI Pooler, which extracts features from candidate bounding box regions based on the overall image features extracted by the backbone network. These extracted features subsequently pass through a self-attention mechanism (SA) and dynamic convolutional layers (DC), which serve to refine the target features further. Temporal information is then utilized to adjust the target features. The refined features are subsequently input into the heads (HD) for classification and regression to determine the category probabilities of objects and predict bounding boxes. This step involves a gradual refinement of the positions and scales of the noisy boxes to converge on the predicted outcomes. In the post-processing module, the predicted results undergo filtering via Non-Maximum Suppression (NMS) and are rescaled to match the original image resolution.
@@
Image encoder. In the design of ConsistencyDet, three distinct backbone architectures have been chosen to function as image encoders: ResNet-50 [25], ResNet-101 [25], and Swin Transformer-base [26]. To facilitate multi-scale representation, which is critical for capturing details across different object sizes, the Feature Pyramid Network [37] is integrated with each of these backbone networks.
@@
Detection decoder. The detection decoder proposed in this study is composed of several cascaded stages that feature specially designed basic modules. Typically, the number of these modules is set to six during the training phase, but this can be adjusted during inference. This architecture inherits its framework from DiffusionDet [21], which is characterized by initiating the detection process with randomly initialized bounding boxes and requiring only those proposal boxes as input. Notably, it employs a detector head that is shared across iterative sampling steps and is orchestrated by time step embeddings [38]. This iterative mechanism distinguishes the detection decoder from those used in other methods, which generally operate in a single-pass fashion, thereby defining both the architecture and the methodology unique to DiffusionDet.
@@
The distinctions between the decoder proposed here and the one used in DiffusionDet can be summarized as follows:
@@
  • The noised boxes input into the detection decoder are not only subject to operations for the addition of Gaussian noise but are also scaled by a coefficient of one-half $(c_{out})$ to constrain the position of the noised boxes. In contrast, DiffusionDet applies a direct clamping operation on the Gaussian-noise-augmented input boxes.
@@
  • ConsistencyDet mandates that the noised boxes at each consecutive pair of time steps $(t-1,t)$ be processed by two distinct detection decoders to generate respective predictions. These predictions are then evaluated against the GT boxes to ensure the model’s consistency. Notably, the noised boxes at the $(t-1)$-th time step are not randomly generated; instead, they are derived through specific computational processes. In contrast, the Diffusion Model solely relies on randomly generated noised boxes at the current $t$-th time step.
3.3 Training
@@
During the training phase, the diffusion process is initialized by creating perturbed bounding boxes derived from the GT boxes. The objective of the model is then to reverse this perturbation, effectively mapping the noisy bounding boxes back to their original GT counterparts. Two distinct training methodologies are employed for ConsistencyDet. The first leverages extant weights from DiffusionDet, incorporating optional knowledge distillation as an enhancement. The second approach is a standalone method that relies on self-training, independent of DiffusionDet. The detailed procedures for each algorithm are delineated in Algorithms 3 and 4. Notably in Algorithm 4, should the knowledge distillation strategy be adopted, the denoiser references the detection outputs from the Diffusion Model. In the absence of knowledge distillation, the variable $x_{s}$ is directly utilized within the denoiser.
@@
       Sample $X_{batch}\in X$; Extract features $E(X_{batch})$; Pad $X_{batch}$ with GT boxes and features as $x_{s}$; Generate a random time step $t_{r}$; // Calculate noise parameters
       Calculate $(\sigma_{t_{r-1}},\sigma_{t_{r}})$ by Eqn. (6); Add noise to $x_{s}$ by Eqn. (7) as $x_{t_{r}}$; // Refer to Algorithm 4
2       Predict $x_{t_{r-1}}$ with $(x_{t_{r}},\sigma_{t_{r}},\sigma_{t_{r-1}},x_{s})$; $d_{t_{r-1}}\leftarrow dcm(x_{t_{r-1}},\sigma_{t_{r-1}})$; $d_{t_{r}}\leftarrow dcm(x_{t_{r}},\sigma_{t_{r}})$; $\mathcal{L}_{t_{r},t_{r}-1}\leftarrow\mathcal{L}(d_{t_{r-1}},G)+\mathcal{L}(d_{t_{r}},G)$; return Loss $\mathcal{L}_{t_{r},t_{r-1}}$
@@
[…]
@@ 4       $x_{0}\leftarrow ddm(x_{t},\sigma_{t})$;
$\nabla_{\sigma}x\leftarrow(x_{t}-x_{0})/\sigma_{t}$; $x_{t-1}\leftarrow x_{t}+\nabla_{\sigma}x\cdot(\sigma_{t-1}-\sigma_{t})$; return $x_{t-1}$
@@
Loss function. The loss functions employed to evaluate the predicted bounding boxes adhere to the framework established by DiffusionDet [21], which incorporates both the $\mathcal{L}_{L1}$ loss and the $\mathcal{L}_{giou}$ loss. The former represents the standard $L1$ loss, while the latter denotes the Generalized Intersection over Union (GIoU) loss as detailed by [41]. Additionally, the classification of each predicted bounding box is assessed using the focal loss, denoted as $\mathcal{L}_{\text{cls}}$. To balance the relative influence of each loss component, a positive real-valued weight $\lambda_{cls/L1/giou}\in\mathbb{R^{+}}$ is allocated to each term. Consequently, the aggregate loss function is formally expressed as:
@@ 3.4 Inference
1 $\Delta t=T/n_{ss}$; Generate random noise $B_{0}$; Extract features $E(X)$; for $t=0$ to $T-1$ step $\Delta t$ do
       Calculate ${\sigma}_{t}$ by Eqn. (6); $x_{0},x_{b},x_{cls},x_{box}\leftarrow P_{c}(E(X),B_{t})$; Perform Box-renewal operation for $x_{b}$ and $x_{0}$; $\nabla_{\sigma}x\leftarrow(x_{b}-x_{0})/{\sigma}_{t}$; $B_{t}\leftarrow x_{b}+\nabla_{\sigma}x(\sigma_{t+\Delta t}-\sigma_{t})$; // Supplement new proposals
@@ 2       $B_{t}\leftarrow conc(B_{t},r([1,n_{p}-n_{r},4])\cdot\sigma_{t+\Delta t})$;
Total time steps. In the simulations conducted, the number of time steps $T$ utilized within the proposed ConsistencyDet framework is configured to 40. This figure represents a notable reduction from the 1000 time steps employed by DiffusionDet, illustrating a marked contrast in the temporal dynamics inherent to the inference processes of the two methodologies. The approach adopted in ConsistencyDet greatly reduces computation time and memory consumption, and may result in only a small decrease in accuracy. During the experiment, it was found that the accuracy did not decrease significantly, but increased in most experiments compared with DiffusionDet.
@@
Box-renewal. Subsequent to each sampling iteration, the predicted outcomes are segregated into two categories: desired and undesired bounding boxes. Desired bounding boxes are those that are precisely aligned with the corresponding objects, whereas undesired bounding boxes remain in positions akin to a random distribution. Propagating these undesired bounding boxes to subsequent iterations is typically counterproductive, as their distributions may quite different with the situations of padded GT boxes with additional noise during the training phase. Therefore, a Box-renewal operation is implemented both in the training and inference phases. This operation entails the discarding of undesired bounding boxes, identified by scores falling beneath a certain threshold. Their subsequent replacement with freshly sampled random boxes, derived from a Gaussian distribution.
4 Experiments

### v5

@@ II-D Nomenclature
III Proposed detection method
In this section, we first introduce the core mathematical principles and underlying theoretical foundations of our method in section III-A. Then, we provide a comprehensive and detailed description of the proposed pipeline in section III-B, which effectively integrates the principles of the Consistency Model into the object detection task. Next, we will delve deeper into a thorough discussion of the specific training strategies and methodologies involved in ConsistencyDet, as shown in section III-C. Finally, we thoroughly examine the complete inference process undertaken by ConsistencyDet in section III-D, including all relevant steps and considerations.
III-A Preliminaries
@@
$p_{\theta}^{(t)}(x_{t-1}|x_{t})=\begin{cases}\mathcal{N}(f_{\theta}^{(1)}(x_{1}),\sigma_{1}^{2}\mathbf{I}),&\text{if }t=1,\\ q_{\sigma}(x_{t-1}|x_{t},f_{\theta}^{(t)}(x_{t})),&\text{otherwise}.\end{cases}$ (2)
@@
Consistency Model. Within the framework of the Consistency Model which utilizes deep neural networks, two cost-effective methodologies are investigated for enforcing boundary conditions. Let $F_{\theta}(x,t)$ represent a free-form deep neural network whose dimensionality is the same as $x$. The first method directly parameterizes the Consistency Model as:
$f_{\theta}(x,t)=\begin{cases}x,&\text{if }t=\tau,\\ F_{\theta}(x,t),&\text{if }t\in[\tau,T),\end{cases}$ (4)
@@
where $\tau$ is an integer in the range $[0,T-1]$. The second method parameterizes the Consistency Model by incorporating skip connections and is formalized as follows:
$f_{\theta}(x,t)=c_{skip}(t)x+c_{out}(t)F_{\theta}(x,t),$ (5)
@@
where $c_{skip}(t)$ and $c_{out}(t)$ are differentiable functions [20], satisfying $c_{skip}(\tau)=1$ and $c_{out}(\tau)=0$. By employing this construction, the Consistency Model becomes differentiable at $t=\tau$, provided that $F_{\theta}(x,t)$, $c_{skip}(t)$, and $c_{out}(t)$ are all differentiable. This differentiability is crucial for the training of continuous-time Consistency Models.
@@
Figure 3: Training procedures of the proposed ConsistencyDet. After extracting features through the backbone, random Gaussian noise is added to GT boxes following the Consistency Model’s noise addition strategy. These noised boxes with corresponding features processed by RoI pooler are then input to ConsistencyHead for iterative noise removal, with several basic modules, ultimately yielding the final detection results. Each the basic module contains a self-attention mechanism (SA), dynamic convolutional layers (DC) and the head (HD) of classification and regression.
@@
Figure 4: To ensure the self-consistency property of the Consistency Model, it requires feeding the noised boxes corresponding to $(t-1)$-th and $t$-th time steps into the model simultaneously. They jointly predict the final results, then compare them with GT to estimate the training loss.
@@
Training process of Consistency Model. Within the context of the Consistency Model, this study introduces two different training strategies that capitalize on the self-consistency property. The proposed strategies encompass distillation training, which transfers knowledge from a Diffusion Model to the Consistency Model, and an autonomous training regime where the Consistency Model is trained independently. The first method employs numerical ODE solvers in conjunction with a pre-trained diffusion-based detector, referred to as DiffusionDet, to generate pairs of proximate points along a PF ODE trajectory. By minimizing the divergence between the model predictions for these point pairs, knowledge from the Diffusion Model is effectively transferred to the Consistency Model. As a result, this allows for the generation of high-quality samples through a single evaluation of the network [31, 32]. The comprehensive training protocol for the Consistency Model through distillation is encapsulated in Algorithm 1. Conversely, the second approach dispenses with the need for a pre-trained Diffusion Model, thereby enabling the Consistency Model to be trained in a standalone manner. This technique delineates the Consistency Model as an independent entity within the broader spectrum of its applications. The detailed training process for the independently trained Consistency Model is encapsulated in Algorithm 2.
@@
1
[…]
3while not convergence do
@@ 9       $\theta^{-}\leftarrow\text{stopgrad}(\mu\theta^{-}+(1-\mu)\theta)$;
2
[…]
4while not convergence do
5       Sample $\mathbf{x}\in\mathcal{D}$, and $t\in\mathcal{U}\left[1,T(k)-1\right]$;
6      Sample $\epsilon\in\mathcal{N}(0,I)$;
7      $\mathcal{L}(\theta,\theta^{-})\leftarrow\lambda(\sigma_{t})d(f_{\theta}\left(\mathbf{x}+\epsilon\sigma_{t+1},\sigma_{t+1}\right),f_{\theta^{-}}\left(\mathbf{x}+\epsilon\sigma_{t},\sigma_{t}\right))$;
8      $\theta\leftarrow\theta-\eta\nabla_{\theta}\mathcal{L}(\theta,\theta^{-})$;
9      $\theta^{-}\leftarrow\text{stopgrad}(\mu(k)\theta^{-}+(1-\mu(k))\theta)$;
10      $k\leftarrow k+1$;
@@ III-B Pipeline
ConsistencyDet is composed of an image encoder and a detection decoder, with the workflow depicted in Fig. 3. The input to the image encoder is the original image along with its GT boxes. These GT boxes are then subjected to a noise injection procedure as part of the Consistency Model, where Gaussian noise is randomly introduced. The perturbed boxes are processed by the RoI Pooler, which extracts features from the regions of candidate bounding boxes based on the overall image features extracted by the backbone network. These extracted features subsequently pass through a self-attention mechanism (SA) and dynamic convolutional layers (DC), which serve to refine the target features further. Temporal information is then utilized to adjust the target features. The refined features are subsequently input into the heads (HD) for classification and regression to determine the category probabilities of objects and predict bounding boxes. This step involves a gradual refinement of the positions and scales of the noisy boxes to converge on the predicted outcomes. In the post-processing module, the predicted results undergo filtering via Non-Maximum Suppression (NMS) and are rescaled to match the original image resolution.
@@
Image encoder. In the design of ConsistencyDet, three different backbone architectures are chosen to function as image encoders: ResNet-50 [23], ResNet-101 [23], and Swin Transformer-base [26]. To facilitate multi-scale representation, which is critical for capturing details across different object sizes, the Feature Pyramid Network [33] is integrated with each of these backbone networks.
@@
Detection decoder. The detection decoder proposed in this study is composed of several cascaded stages that feature specially designed basic modules. Typically, the number of these modules is set to six during the training phase, but this can be adjusted during inference. This architecture inherits its framework from DiffusionDet [19], which is characterized by initiating the detection process with randomly initialized bounding boxes and requiring only those proposal boxes as input. Notably, it employs a detector head that is shared across iterative sampling steps and is orchestrated by time step embeddings [34]. This iterative mechanism distinguishes the detection decoder from those used in other methods, which generally operate in a single-pass fashion, thereby defining both the architecture and the methodology unique to ConsistencyDet.
@@
The differences between the decoder proposed here and the one used in DiffusionDet can be summarized as follows:
@@
  • The noised boxes fed into the detection decoder are not only subject to operations for the addition of Gaussian noise but are also scaled by a coefficient, $c_{in}(t)/2$, to constrain the position of the noised boxes. In contrast, DiffusionDet applies a direct clamping operation on the Gaussian-noise-augmented input boxes.
@@
  • ConsistencyDet mandates that the noised boxes at each consecutive pair of time steps $(t\!\!-\!\!1,t)$ be processed by two different detection decoders to generate respective predictions. These predictions are then evaluated against the GT boxes to ensure the model’s consistency. Notably, the noised boxes at the $(t\!\!-\!\!1)$-th time step are not randomly generated; instead, they are derived through specific computational processes. In contrast, the Diffusion Model solely relies on randomly generated noised boxes at the current $t$-th time step.
III-C Training
@@
During the training phase, the diffusion process is initialized by creating perturbed bounding boxes derived from the GT boxes. The objective of the model is then to reverse this perturbation, effectively mapping the noisy bounding boxes back to their original GT counterparts. Two different training methodologies are employed for ConsistencyDet. The first leverages extant weights from DiffusionDet, incorporating optional knowledge distillation as an enhancement. The second approach is a standalone method that relies on self-training, independent of DiffusionDet. The detailed procedures for each algorithm are delineated in Algorithms 3 and 4. Notably in Algorithm 4, should the knowledge distillation strategy be adopted, the denoiser references the detection outputs from the Diffusion Model. In the absence of knowledge distillation, the variable $x_{s}$ is directly utilized within the denoiser.
1
@@
3       Sample $X_{batch}\in X$;
4      Extract features $E(X_{batch})$;
5      Pad $X_{batch}$ with GT boxes and features as $x_{s}$;
6      Generate a random time step $t_{r}$;
      // Calculate noise parameters
7
8      Calculate $(\sigma_{t_{r-1}},\sigma_{t_{r}})$ by Eqn. (6);
9      Add noise to $x_{s}$ by Eqn. (7) as $x_{t_{r}}$;
      // Refer to Algorithm 4
10
11      Predict $x_{t_{r-1}}$ with $(x_{t_{r}},\sigma_{t_{r}},\sigma_{t_{r-1}},x_{s})$;
12      $d_{t_{r-1}}\leftarrow dcm(x_{t_{r-1}},\sigma_{t_{r-1}})$;
13      $d_{t_{r}}\leftarrow dcm(x_{t_{r}},\sigma_{t_{r}})$;
14      $\mathcal{L}_{t_{r},t_{r}-1}\leftarrow\mathcal{L}(d_{t_{r-1}},G)+\mathcal{L}(d_{t_{r}},G)$;
15      return Loss $\mathcal{L}_{t_{r},t_{r-1}}$
16
@@ 14      $\mathcal{L}_{t_{r},t_{r}-1}\leftarrow\mathcal{L}(d_{t_{r-1}},G)+\mathcal{L}(d_{t_{r}},G)$;
GT boxes padding. In open-source benchmarks for object detection, such as those referenced in [35, 25, 36], there is typically a variance in the number of annotated instances across different images. To address this inconsistency, we implement a padding strategy that introduces auxiliary boxes around the GT boxes. This ensures that the total count of boxes attains a predetermined number, $n_{tr}$, during the training phase. These padded instances are denoted as $x_{s}$, signifying the original padded samples. For the $i$-th GT box, represented by $\textit{b}^{\textit{i}}$, we apply Gaussian noise to its four parameters $(\textit{c}_{x}^{\textit{i}},\textit{c}_{y}^{\textit{i}},\textit{w}^{\textit{i}},\textit{h}^{\textit{i}})$ at a randomly selected time step $t$.
1
@@ 6       $x_{0}\leftarrow ddm(x_{t},\sigma_{t})$;
[…]
10return $x_{t-1}$
11
@@
Loss function. The loss functions employed to evaluate the predicted bounding boxes adhere to the framework established by DiffusionDet [19], which incorporates both the $\mathcal{L}_{L1}$ loss and the $\mathcal{L}_{giou}$ loss. The former represents the standard $L1$ loss, while the latter denotes the Generalized Intersection over Union (GIoU) loss as detailed by [37]. Additionally, the classification of each predicted bounding box is assessed using the focal loss, denoted as $\mathcal{L}_{cls}$. To balance the relative influence of each loss component, a positive real-valued weight $\lambda_{cls/L1/giou}\in\mathbb{R^{+}}$ is allocated to each term. Consequently, the aggregate loss function is formally expressed as:
@@ III-D Inference
1
[…]
3Generate random noise $B_{0}$;
4Extract features $E(X)$;
5for $t=0$ to $T-1$ step $\Delta t$ do
6       Calculate ${\sigma}_{t}$ by Eqn. (6);
7      $x_{0},x_{b},x_{cls},x_{box}\leftarrow P_{c}(E(X),B_{t})$;
8      Perform Box-renewal operation for $x_{b}$ and $x_{0}$;
9      $\nabla_{\sigma}x\leftarrow(x_{b}-x_{0})/{\sigma}_{t}$;
10      $B_{t}\leftarrow x_{b}+\nabla_{\sigma}x(\sigma_{t+\Delta t}-\sigma_{t})$;
      // Supplement new proposals
11
@@ 12      $B_{t}\leftarrow conc(B_{t},r([1,n_{p}-n_{r},4])\cdot\sigma_{t+\Delta t})$;
Total time steps. In the simulations conducted, the number of time steps $T$ utilized within the proposed ConsistencyDet framework is configured to 40. This figure represents a notable reduction from the 1000 time steps employed by DiffusionDet, illustrating a marked contrast in the temporal dynamics inherent in the inference processes of the two methodologies. The approach adopted in ConsistencyDet greatly reduces computation time and memory consumption, and may result in only a small decrease in accuracy. During the experiment, it was found that accuracy did not decrease significantly, but increased in most experiments compared with DiffusionDet.
@@
Box-renewal. Subsequent to each sampling iteration, the predicted outcomes are segregated into two categories: desired and undesired bounding boxes. Desired bounding boxes are those that are precisely aligned with the corresponding objects, whereas undesired bounding boxes remain in positions akin to a random distribution. Propagating these undesired bounding boxes to subsequent iterations is typically counterproductive, as their distributions may be quite different with the situations of padded GT boxes with additional noise during the training phase. Therefore, a Box-renewal operation is implemented both in the training and inference phases. This operation entails the discarding of undesired bounding boxes, identified by scores falling beneath a certain threshold. Their subsequent replacement with freshly sampled random boxes, is derived from a Gaussian distribution.
IV Experiment